OpenStudy (anonymous):

find the first three nontrivial terms in the power series expansions centered at zero for the following functions: a) f(x) = e^(-2x) sin(x) b) g(x) = x^2 3^x + sin(x+2)

OpenStudy (anonymous):

someone help me please

OpenStudy (anonymous):

\(\sin(x)=x-\frac{x^3}{3!}+\frac{x^5}{5!}-...\)

OpenStudy (anonymous):

\[e^{-2x}=1-2 x+2 x^2-\frac{4 x^3}{3}+...\]

OpenStudy (anonymous):

multiply the first few terms and see what you get

OpenStudy (anonymous):

So you would get = x + 2x^5/3! + 2x^7/5! etc..... How did you get the series for sin and e?

OpenStudy (anonymous):

i memorized the one for sine it is a very common function, good to just know it here is an easy way to remember it: \(\sin(0)=0\) so there is no constant term and sine is odd, so only odd powers and it alternates

OpenStudy (anonymous):

also memorize the one for \(e^x\) which is very easy \[\sum_{n=0}^{\infty}\frac{x^n}{n!}\] then replace \(x\) by \(-2x\)

OpenStudy (anonymous):

but i am fairly sure you did not multiply correctly

OpenStudy (anonymous):

you have \(1\times x\) as one term next lowest power comes from \(-2x\times x=-2x^2\) then the cube term comes from \(-\frac{x^3}{3!}\times 1\)

OpenStudy (anonymous):

so i think it begins \[x-2x^2-\frac{x^3}{3}+...\]